An electric motor accelerates a grinding wheel at constant acceleration from 60 rev/min to 287 rev/min in 11 seconds. How many revolutions did it make in this time? Answer in units of rev
Answer #1

Solution)

Here,

increase in angular speed = 287 - 60 = 227 rpm

Time, t = 11 s

Increase in angular speed= w = 2π(227/60) = 23.75 rad/s

α = w/t = 23.75/11 = 2.15 rad/s²

Now,

Θ = (0.5)(2.15)(11)² = 130.67 rad

So,

No. of rotations of grinding wheel = 130.6/2π ≈ 20.80

==========